OpenStudy (anonymous):

Let f(x)=x^(1/2). Form and simplify the difference quotient [f(x+h) + f(x)]/h.

OpenStudy (turingtest):

these problems are almost always about rationalizing the fraction

OpenStudy (turingtest):

i.e. whenever you are trying to use the difference quotient on an expression under a radical it pays to try this:\[{f(x-h)-f(x)\over h}={\sqrt{x-h}-\sqrt x\over h}={\sqrt{x-h}-\sqrt x\over h}\cdot{\sqrt{x-h}+\sqrt x\over \sqrt{x-h}+\sqrt x}\]

OpenStudy (anonymous):

Would I multiply both the numerator and the denominator by 1/(x+h)^1/2 +1/x^1/2 or by (x+h)^1/2 +x^1/2

OpenStudy (turingtest):

*sorry that should all be +h in the above a contagious typo from your post :P\[{f(x+h)-f(x)\over h}={\sqrt{x+h}-\sqrt x\over h}={\sqrt{x+h}-\sqrt x\over h}\cdot{\sqrt{x+h}+\sqrt x\over \sqrt{x+h}+\sqrt x}\]

OpenStudy (turingtest):

you multiply by the conjugate of the numerator the conjugate of a radical\[\sqrt a\pm\sqrt b\]is\[\sqrt a\mp\sqrt b\]

OpenStudy (turingtest):

so because in the top we have \(\sqrt{x+h}-\sqrt x\) we need to multiply it by the its conjugate, which is \(\sqrt{x+h}+\sqrt x\) of course, that means we need to multiply the bottom by that as well this trick is called "rationalization"

OpenStudy (turingtest):

what do you get in the numerator after multiplying by \(\sqrt{x+h}+\sqrt x\) ?

OpenStudy (anonymous):

1/(x+h) - 1/x

OpenStudy (turingtest):

just do the numerator\[(\sqrt{x+h}-\sqrt x)(\sqrt{x+h}+\sqrt x)=?\]

OpenStudy (anonymous):

\[(x+h) - h\]

OpenStudy (anonymous):

er I'm sorry (x+h) - x which would jsut be h

OpenStudy (turingtest):

correct, which simplifies to just \(h\) so \(h\) is in the numerator, and in the denominator is what?

OpenStudy (anonymous):

\[h \sqrt{x+h} + h \sqrt{x} \]

OpenStudy (turingtest):

yes so can we cancel anything?

OpenStudy (anonymous):

factor out the h

OpenStudy (turingtest):

well, cancel out the h we got one on top and bottom so it's gone

OpenStudy (anonymous):

\[1/\sqrt{x+h} + \sqrt{x}\]

OpenStudy (turingtest):

yeah, which we can take the limit of now :)

OpenStudy (turingtest):

you should really have put parentheses around the denominator, but I know what you mean

OpenStudy (anonymous):

Yeah sorry, first time using this site, still getting use to the equation maker

OpenStudy (turingtest):

it's all good, and welcome to OS :) so the limit of the above as h goes to zero is...?

OpenStudy (anonymous):

\[1/(2 \sqrt{x})\]

OpenStudy (turingtest):

right, which if you know the power rule you know is the correct answer

OpenStudy (anonymous):

Haha, yeah, the power rule makes everything easier. Lots of arithmetic on that problem. Would that count as the simplified version?

OpenStudy (turingtest):

the power rule has to be proven... using it is like a shortcut, so when they ask for the difference quotient you should do it if you have to show your work you can always check your final answer with the power rule though

OpenStudy (anonymous):

Awesome, thanks very much!

OpenStudy (turingtest):

welcome :)